OpenStudy (anonymous):

Anna has been studying a type of bacteria that triples every month. Originally, there were 4 bacterial cells. She wants to know how many there will be after 15 months. Which equation should she use

OpenStudy (anonymous):

a)a15=4(3)(15) b)a15=4(3)^15 c)a15=4(3)^15-1 d)a15-4(3)(15-1)

OpenStudy (anonymous):

4 (3) \[^{15}\]\[4 * 3^{15}\]